A goldsmith has two gold alloys. The first alloy is 30% gold; the second alloy is 80% gold. How many grams of each should be mixed to produce 40 grams of an alloy that is 70% gold?
Amount of 30% gold in grams?
Amount of 80% gold in grams??

Answer by ankor@dixie-net.com(22740)

A goldsmith has two gold alloys.
The first alloy is 30% gold; the second alloy is 80% gold.
How many grams of each should be mixed to produce 40 grams of an alloy
that is 70% gold?
:
Let x = amt of 80% alloy required
the resulting total is to be 40 grams, therefore:
(40-x) = amt of 30% gold
:
.80x + .30(40-x) = .70(40)
.80x + 12 - .30x = 28
.80x - .30x = 28 - 12
.50x = 16
x = 16/.5
x = 32 grams of the 80% alloy
then
40 - 32 = 8 grams of 30% alloy